### v1

@@ 1 Introduction
Figure 1 presents the results of an AL experiment on CIFAR-10 with a total budget of 210 instances and a batch acquisition size of 10 instances per cycle. The detailed experimental setup is available in the Appendix LABEL:app:graphical_abstract. We evaluate prominent state-of-the-art AL strategies, focusing on performance (i.e., accuracy) and acquisition times (i.e., required time for selecting instances per cycle). Fig. 1(a) plots the accuracy difference between a model trained with an AL strategy and a model trained with randomly sampled instances. Notably, Bait outperforms every strategy, demonstrating its superior performance. However, the acquisition times in Fig. 1(b) indicate that Bait requires significantly more time than others to acquire a batch of instances. This bottleneck will become even more severe when dealing with larger acquisition sizes or more classes. Consequently, despite Bait’s superior accuracy, this example highlights its computational challenges. These practical constraints limit its feasibility across various applications and real-world settings.
@@ 6.1 Setup
Datasets: We conduct experiments across a variety of image datasets to assess the effectiveness and robustness of our proposed approximations of Bait. Our evaluation encompasses a total of nine image datasets. Aiming to ensure a comprehensive evaluation focusing on scalability, we cover a broad range of class numbers. These datasets range from a small (CIFAR-10 [20], STL-10 [39], and Snacks [29]) to a large amount of classes (CIFAR-100 [20], Food-101 [6], Flowers-102 [30], StanfordDogs [42], TinyImageNet [22], ImageNet [10]). For each dataset, we use a 10% split of the training dataset for development of our approximations. Information about the number of instances and classes is summarized in Tab. 2. A more comprehensive description can be found in Appendix LABEL:app:datasets.
@@
AL Setting: We determine the acquisition size, total budget, and the number of initial instances according to the dataset characteristics to guarantee convergence of the AL process. Additionally, the initial labeled pool is chosen randomly. Details are provided in Appendix LABEL:app:al_settings. Our evaluation comprises various AL selection strategies, including Random, which selects instances randomly, Margin [5], focusing on the uncertainty in the top-2 predicted probabilities, Badge [4], which selects instances according to the gradient norm of the likelihood, and Typiclust [15], employing clustering and density estimation for selection. These strategies are considered state-of-the-art and were chosen based on their robustness and efficiency across different domains [5, 33, 15]. Additionally, we employ Bait with our expectation approximation, called Bait (Exp), and our binary approximation, called Bait (Binary).
Metrics: To evaluate the effectiveness of a strategy, we consider its learning curve. Therefore, in each cycle, we train a DNN on a labeled pool selected by a given strategy and compute the accuracy on the test dataset.Then, we examine the accuracy improvement over Random by considering the difference of learning curves. Due to space limitations, we only report a selection of learning curves. All remaining curves of all strategies and datasets can be found in Appendix LABEL:app:learning_curves. As an alternative, we report the area under the learning curve (AUC), which summarizes the learning curve into a single numerical value. All learning curves and numerical values are averaged over ten repetitions to ensure comparability.
@@ 6.2 Assessment of Approximations
Bait (Binary): We examine the effectiveness of our second approximation discussed in Section 5.2, which reformulates the multi-class classification task into a binary one. Since this approximation is intended to work on a large number of classes, we also compare it to the diagonal approximation of the FIM. This approximation only calculates the main diagonal, omitting off-diagonal elements. It is frequently employed to scale the FIM to high dimensions [23], and is therefore suitable for this comparison. We refer to Appendix LABEL:app:diagonal_fim for a more detailed explanation. Considering Tab. 4, we notice that the binary approximation yield similar (or better) accuracies compared to the original version of Bait across all datasets. In contrast, the diagonal FIM approximation is not able to achieve comparable performances. Regarding the time complexity, we see that both the diagonal and binary approximation of Bait significantly reduce the time complexity. While the diagonal approximation is slightly slower, our approximation is able to acquire instances quicker and independent of the class number. This is noticeable due to (almost) constant acquisition times between Snacks and STL-10, which both have an identical number of instances but a different number of classes. To ensure that not only the AUC values but also the behavior of the approximations are the similar, we show accuracy improvement curves in Fig. 2 for Bait and its approximations. They show that our approximations Bait (Exp) and Bait (Binary) closely resemble the original formulation, while the diagonal approximation shows strong divergence.
@@ 6.3 Benchmark Experiments
Next, we present the main results on all datasets to compare the approximations of Bait to other state-of-the-art AL strategies. Table 5 shows the average AUC improvement over Random of all strategies across all datasets. In the case of Bait (Exp), memory limitations make scaling beyond 50 classes infeasible, thus we only report results on the three datasets with less than 100 classes. The learning curves and accuracy improvement curves for each dataset can be found in Appendix LABEL:app:learning_curves.

### v2

@@ 1 Introduction
Figure 1 presents the results of an AL experiment on CIFAR-10 with a total budget of 210 instances and a batch acquisition size of 10 instances per cycle. The detailed experimental setup is available in the Appendix 0.A. We evaluate prominent state-of-the-art AL strategies, focusing on performance (i.e., accuracy) and acquisition times (i.e., required time for selecting instances per cycle). Fig. 1(a) plots the accuracy difference between a model trained with an AL strategy and a model trained with randomly sampled instances. Notably, Bait outperforms every strategy, demonstrating its superior performance. However, the acquisition times in Fig. 1(b) indicate that Bait requires significantly more time than others to acquire a batch of instances. This bottleneck will become even more severe when dealing with larger acquisition sizes or more classes. Consequently, despite Bait’s superior accuracy, this example highlights its computational challenges. These practical constraints limit its feasibility across various applications and real-world settings.
@@ 6.1 Setup
Datasets: We conduct experiments across a variety of image datasets to assess the effectiveness and robustness of our proposed approximations of Bait. Our evaluation encompasses a total of nine image datasets. Aiming to ensure a comprehensive evaluation focusing on scalability, we cover a broad range of class numbers. These datasets range from a small (CIFAR-10 [20], STL-10 [39], and Snacks [29]) to a large amount of classes (CIFAR-100 [20], Food-101 [6], Flowers-102 [30], StanfordDogs [42], TinyImageNet [22], ImageNet [10]). For each dataset, we use a 10% split of the training dataset for development of our approximations. Information about the number of instances and classes is summarized in Tab. 2. A more comprehensive description can be found in Appendix 0.B.
@@
AL Setting: We determine the acquisition size, total budget, and the number of initial instances according to the dataset characteristics to guarantee convergence of the AL process. Additionally, the initial labeled pool is chosen randomly. Details are provided in Appendix 0.D. Our evaluation comprises various AL selection strategies, including Random, which selects instances randomly, Margin [5], focusing on the uncertainty in the top-2 predicted probabilities, Badge [4], which selects instances according to the gradient norm of the likelihood, and Typiclust [15], employing clustering and density estimation for selection. These strategies are considered state-of-the-art and were chosen based on their robustness and efficiency across different domains [5, 33, 15]. Additionally, we employ Bait with our expectation approximation, called Bait (Exp), and our binary approximation, called Bait (Binary).
Metrics: To evaluate the effectiveness of a strategy, we consider its learning curve. Therefore, in each cycle, we train a DNN on a labeled pool selected by a given strategy and compute the accuracy on the test dataset. Then, we examine the accuracy improvement over Random by considering the difference of learning curves. Due to space limitations, we only report a selection of learning curves. All remaining curves of all strategies and datasets can be found in Appendix 0.E. As an alternative, we report the area under the learning curve (AUC), which summarizes the learning curve into a single numerical value. All learning curves and numerical values are averaged over ten repetitions to ensure comparability.
@@ 6.2 Assessment of Approximations
Bait (Binary): We examine the effectiveness of our second approximation discussed in Section 5.2, which reformulates the multi-class classification task into a binary one. Since this approximation is intended to work on a large number of classes, we also compare it to the diagonal approximation of the FIM. This approximation only calculates the main diagonal, omitting off-diagonal elements. It is frequently employed to scale the FIM to high dimensions [23], and is therefore suitable for this comparison. We refer to Appendix 0.C for a more detailed explanation. Considering Tab. 4, we notice that the binary approximation yield similar (or better) accuracies compared to the original version of Bait across all datasets. In contrast, the diagonal FIM approximation is not able to achieve comparable performances. Regarding the time complexity, we see that both the diagonal and binary approximation of Bait significantly reduce the time complexity. While the diagonal approximation is slightly slower, our approximation is able to acquire instances quicker and independent of the class number. This is noticeable due to (almost) constant acquisition times between Snacks and STL-10, which both have an identical number of instances but a different number of classes. To ensure that not only the AUC values but also the behavior of the approximations are the similar, we show accuracy improvement curves in Fig. 2 for Bait and its approximations. They show that our approximations Bait (Exp) and Bait (Binary) closely resemble the original formulation, while the diagonal approximation shows strong divergence.
@@ 6.3 Benchmark Experiments
Next, we present the main results on all datasets to compare the approximations of Bait to other state-of-the-art AL strategies. Table 5 shows the average AUC improvement over Random of all strategies across all datasets. In the case of Bait (Exp), memory limitations make scaling beyond 50 classes infeasible, thus we only report results on the three datasets with less than 100 classes. The learning curves and accuracy improvement curves for each dataset can be found in Appendix 0.E.
@@ 7 Conclusion
Appendix 0.A Experimental Setup
The graphical abstract is based on the same experimental setup as the main experiments. For the Active Learning (AL) setting, we use the CIFAR-10 dataset consisting of ten classes. We start with 0 labeled instances and select 10 instances per cycle for a total of 21 cycles. In the first cycle, all instances are sampled randomly. We employ the DINOv2-ViT-S/14 model [31] with a fully connected linear classification layer. In each AL cycle, we randomly initialize the classification head for 200 epochs and employ the Rectified Adam optimizer [26]. We use a training batch size of 128, a learning rate of 0.2, and weight decay of 0.0001. In addition, we utilize a cosine annealing learning rate scheduler.
Appendix 0.B Dataset Description
Here, we describe the datasets used in the main experiments in more detail. CIFAR [20] consists of 60,000 colored images with a low $32\times 32$ resolution. There is a predetermined split of 50,000 images as training samples and 10,000 images as test samples. Two variants of this dataset exist, distinguished by their differing number of classes. CIFAR-10 is a coarse-grained task with broad classes such as automobile, dog, airplane, and ship. The classes are mutually exclusive meaning there is no overlap between, e.g., automobiles and trucks. CIFAR-100, containing 100 classes, provides more fine-grained differentiation, such as bicycle, apple, television, and roses. Food-101 [6] consists of 101 food categories with 750 training and 250 test images per category, making a total of 101,000 images. Snacks is a smaller dataset, containing 20 different classes of snack foods, and is split into 5,000 training images and 1,000 test images. Flowers-102 [30] consists of 102 flower categories, commonly occurring in the United Kingdom. Each class consists of 40 to 258 images, making it an imbalanced dataset with a total of 8,189 images. The images have large scale, pose, and light variations. In addition, some categories have large variations within the category and several very similar categories. StanfordDogs [42] contains 20,580 images of 120 classes of dogs from around the world. It is partitioned into 12,000 images for training and 8,580 images for testing. ImageNet [10] is a large-scale image database with 14,197,122 labeled images. They are organized by the semantic hierarchy of WordNet. Images vary in size and are, therefore, usually cropped to 224×224. A common subset is the ImageNet1k dataset, containing 1,281,167 training images, 100,000 test images, and an additional 50,000 validation images, which can be classified into 1,000 classes. In our work, we refer to ImageNet1k as ImageNet. Tiny ImageNet [22] is a scaled-down subset of ImageNet. It consists of 200 classes, and for each class, it provides 500 training images, 50 validation images, and 50 test images. All images have been downsampled to 64×64. STL-10 [39] is another derivation of ImageNet commonly used to evaluate algorithms of unsupervised feature learning. It contains 100,000 unlabeled images and 13,000 labeled images. The labeled images are divided into 5,000 training images and 8,000 test images. All images have been downsampled to 96×96.
Appendix 0.C Diagonal Approximation of the Fisher Information Matrix
A common approximation of the Fisher information matrix (FIM) is to only consider the diagonal elements [23]. This simplifies the outer product considerably reducing the time and space complexity:
$\boldsymbol{I}(\boldsymbol{x};\boldsymbol{\theta})\approx\mathbb{E}_{y}\left[\operatorname{diag}(\nabla_{\boldsymbol{\theta}}\ln p({y}|\boldsymbol{x},\boldsymbol{\theta})\odot\nabla_{\boldsymbol{\theta}}\ln p({y}|\boldsymbol{x},\boldsymbol{\theta}))\right]$ (7)
where $\odot$ is the Hadamard product. This basically omits the off-diagonal elements, which represent the correlations between parameters. This can lead to a less accurate understanding of the parameter space, especially in models where parameter interactions are significant.
Table 6: Overview of AL-related hyperparameters for each dataset.
Dataset Initial Pool Size Acquisition Size Number of Acquisitions
CIFAR10 10 10 20
CIFAR100 100 100 20
Food-101 100 100 20
Flowers-102 50 50 20
StanfordDogs 100 100 20
Tiny ImageNet 200 200 20
ImageNet 1000 1000 20
Appendix 0.D Active Learning Settings
In Tab. 6, we list AL related settings. We derived these values by qualitatively investigating the learning curves of Random sampling in initial experiments with a 10% split of the training dataset. This way, we try to ensure convergence of learning curves across strategies, which is necessary for the area under the learning curve of a strategy to be representative.
Appendix 0.E Learning Curves
All learning curves from the main experiments are depicted in Fig. 45, and 6. For more details, we refer to our implementation.
Accuracy
Number of Annotations
Accuracy Difference to Random
Number of AnnotationsCIFAR-10
Accuracy
Number of Annotations
Accuracy Difference to Random
Number of AnnotationsSTL-10
Accuracy
Number of Annotations
Accuracy Difference to Random
Number of AnnotationsSnacks
Accuracy
Number of Annotations
Accuracy Difference to Random
Number of AnnotationsCIFAR-100
Accuracy
Number of Annotations
Accuracy Difference to Random
Number of AnnotationsFood-101
Accuracy
Number of Annotations
Accuracy Difference to Random
Number of AnnotationsFlowers-102
Accuracy
Number of Annotations
Accuracy Difference to Random
Number of AnnotationsStandfordDogs
Accuracy
Number of Annotations
Accuracy Difference to Random
Number of AnnotationsTinyImagenet
Accuracy
Number of Annotations
Accuracy Difference to Random
Number of AnnotationsImagenet